The energy of a mole of photons of light in the electromagnetic spectrum decreases logarithmically as a function of increasing wavelength. In the visible range, light has between 40 and 80 kcal/mole of photon energy. Below about 400 nm wavelength, the energy increases significantly. Use what you know about the energy of chemical bonds to make an argument in favor of using sun screen with UV-block on a sunny day.
Answer #1

energy and wavwlength are related by the formula

E = hc/L

E is energy

h is palcnks constant

c is velocity of light

L is wavelength

so
for 400nm wavelength

Energy E = (6.626 *10^-34 * 3*10^8)/(400 *10^-9)

E = 4.9*10^-19 Joules

or

E = 4.9 *10^-19/(1.6*10^-19)

E = 3.1 eV


also 1 eV = 23 K cal/mole

so

3.1 eV = 71.49 K cal/mole